# Power-Law Running of the Higgs Mass

### 16 Mar 2026 v1

@@ Abstract
The renormalized scalar mass squared is a function of the external momentum $q^{2}$ and power-runs as $q^{2}$: it is $\mathcal{O}(M_{\text{GUT}}^{2})$ at $M_{\text{GUT}}^{2}$ and $\mathcal{O}(m_{W}^{2})$ at $m_{W}^{2}$. There is no hierarchy problem.
@@
The fully renormalized mass function virtual ; Choi:2023:observables ; Choi:2025:Higgs shows that the scalar mass squared runs as $q^{2}$. This is not an additional assumption but a consequence of the standard quantum field theory tHooft:1971 ; tHooft:Veltman:1972 , once the renormalization is carried out completely. Therefore it is $\mathcal{O}(M_{\text{GUT}}^{2})$ at $M_{\text{GUT}}^{2}$ and $\mathcal{O}(m_{W}^{2})$ at $m_{W}^{2}$—the natural scaling of an unprotected quantity. In a grand unified theory, an order-one boundary condition at the unification scale suffices to produce an electroweak-scale Higgs, replacing the traditional $10^{-28}$ cancellation.
@@
The mass function constructed in Ref. virtual defines the scalar mass at every momentum. On-shell renormalization yields Choi:2023:observables ; Choi:2025:Higgs
@@
with the dressed propagator $i/(q^{2}-m^{2}(q^{2}))$ and pole mass $m_{h}$ fixed by $m^{2}(m_{h}^{2})=m_{h}^{2}$.
The mass function is finite Bogoliubov:Parasiuk:1957 ; Hepp:1966 ; Zimmermann:1969 , and regularization-independent Choi:2023:observables ; Choi:2024:decoupling . In particular, the $q^{2}$ dependence of $m^{2}(q^{2})$ is a physical effect that is present regardless of the renormalization scheme. The contribution from gauge bosons can be made gauge-parameter invariant by absorbing non-invariant terms to the vertices, thanks to the Ward–Takahashi identity. The result coincides with the Feynman gauge choice, which we assume from now on Cornwall:1982 ; Binosi:Papavassiliou:2009 ; virtual .
What runs is the mass function $m^{2}(q^{2})$—the full momentum-dependent dressing. The pole mass is a fixed number; it does not run. The power-law behavior at momenta far from the pole is a prediction; the pole mass itself is data. We should not attribute to the pole mass properties that belong to the mass function.
@@
    This verifies the ’t Hooft account (1). The crucial difference is that the ’t Hooft estimate gives a correction at $M_{\text{GUT}}$, whereas the renormalized mass function $m^{2}(q^{2})$ is defined at every scale and is computed bottom-up from the measured pole mass. Using $m_{h}=125.08$ GeV ATLAS:2023 ; CMS:2025 , the complete one-loop correction to the Higgs mass parameter (Eqs. (26)–(29) of Ref. Choi:2025:Higgs ) gives, at $\sqrt{q^{2}}=M_{\text{GUT}}=10^{16}$ GeV,
@@
    which is $\mathcal{O}(1)$. The mass function, measured in units of the scale at which it is probed, is order one at every scale. The 28 orders of magnitude between $m_{h}^{2}$ and $M_{\text{GUT}}^{2}$ are the natural scaling of a dimension-2 quantity that runs as $q^{2}$—not a hierarchy that requires explanation. The reason is nothing but the absence of symmetry protection of the scalar mass.
@@
In the $\overline{\text{MS}}$ scheme Bardeen:Buras:Duke:Muta:1978 with dimensional regularization, quadratic powers of the renormalization scale $\mu^{2}$ are never generated—they correspond to poles at $d=2$, which are invisible at $d=4-\epsilon$. The beta function for $\bar{m}^{2}(\mu)$ is therefore purely logarithmic, and the scalar mass appears to run like a fermion mass. With no $\mu^{2}$ term in sight, there is nothing to trigger alarm—and the power-law $q^{2}$ dependence, which lives in the finite parts of the amplitude, goes unrecognized. This does not mean that dimensional regularization is at fault—the mass function $m^{2}(q^{2})$ is regularization-independent, as noted above—but that $\overline{\text{MS}}$ is an unfortunate scheme for exposing the physical scale dependence.
The validity of these results requires that perturbation theory remain under control at all scales up to $M_{\text{GUT}}$ Degrassi:2012 ; Buttazzo:2013 . The one-loop correction $\Sigma_{\text{ren}}^{\text{SM}}/q^{2}\approx 0.86$ at $M_{\text{GUT}}$ is less than unity, so the propagator denominator $q^{2}-m^{2}(q^{2})=q^{2}(1-0.86)>0$ remains positive—no sign change, no ghost, no breakdown. The top quark alone gives +1.135+1.135, which exceeds unity, but the $W$ and $Z$ contributions (0.28-0.28 combined) bring the total below one; the full SM content is essential (see Appendix). This partial cancellation is not accidental: fermion loops carry a factor (1)(-1) from the closed Grassmann trace (spin-statistics theorem Pauli:1940 ), giving the top a positive contribution while the $W$ and $Z$ contribute with opposite sign. We have verified that the sign structure persists at two loops—each fermion loop carries its (1)(-1)—so the partial cancellation is a structural property of the SM’s spin content, not a one-loop coincidence. The two-loop correction is suppressed by a further factor of $m_{\text{EW}}^{2}/(16\pi^{2}v^{2})\log(M_{\text{GUT}}^{2}/m_{\text{EW}}^{2})$, giving $\Sigma^{(2)}/\Sigma^{(1)}\sim 0.1$$0.2$; this shifts the total from $0.86$ to roughly 0.90.91.01.0, which does not change the order-one character of $m^{2}(M_{\text{GUT}}^{2})/M_{\text{GUT}}^{2}$. That $r$ approaches unity at $M_{\text{GUT}}$ is expected: this is the boundary of the SM’s perturbative domain, where the GUT description takes over. Perturbation theory is controlled at every scale from the electroweak to the GUT.
@@
We now examine what this implies for a UV completion. The Standard Model must be completed at some high scale, whether by grand unification, string theory, or another framework. The scale $M_{\text{GUT}}\sim 10^{16}$ GeV indicated by gauge coupling unification Georgi:Quinn:Weinberg:1974 ; Langacker:Luo:1991 is representative of where such new physics is expected to appear. We take $SU(5)$ grand unification Georgi:Glashow:1974 as a concrete prototype; the specific parameters differ between UV completions, but the essential structure—an order-one boundary condition at the high scale, with the SM running below it—does not.
@@
where $\mu^{2}$ is the $\mathbf{5}$-Higgs mass parameter, $\lambda_{5}$ the coupling to the adjoint, $V$ the adjoint VEV, and $\Sigma^{\text{heavy}}_{\text{GUT}}$ contains loops from $X$, $Y$ bosons, the triplet, and the adjoint. This is the matching and provides a constraint on the GUT.
@@
an order-one requirement on $\mathcal{O}(1)$ parameters. In the traditional framework, the same constraint was $\sim 10^{-28}$—a cancellation to 28 decimal places. The target has widened to a band of width $\sim 0.3$. The specific value $m_{h}=125$ GeV within the electroweak band is a precise constraint on the GUT—analogous to the constraint that the top Yukawa is $0.99$ rather than 0.50.5—but it is a constraint on $\mathcal{O}(1)$ parameters, not a fine-tuning.
@@
To sum up, the power-law running makes $m^{2}(q^{2})$ scale as $q^{2}$, so $m^{2}(q^{2})/q^{2}$ is $\mathcal{O}(1)$ at every scale. Any order-one GUT-scale boundary condition yields an electroweak-scale pole mass through the SM running. If the scalar mass ran logarithmically like a fermion mass, $m^{2}$ would barely change between scales, and $(125~\text{GeV})^{2}$ from $\mathcal{O}(M_{\text{GUT}}^{2})$ inputs would require $10^{-28}$—the very fine-tuning that the hierarchy problem describes.
@@
Appendix A Explicit one-loop formulas
@@ Acknowledgements.
The author is grateful to Jaehoon Jeong, Taehyun Jung, Hyun-Min Lee, Matthew McCollough and Minho Son for discussions. This work is partly supported by the grant RS-2023-00277184 of the National Research Foundation of Korea. The author used Claude (Anthropic) for editing and discussion during the manuscript preparation. The final content was reviewed and approved by the author, who takes full responsibility for the work.

### 18 Mar 2026 v2

@@ Abstract
The renormalized scalar mass squared is a function of the energy scale and power-runs as its square: it is $\mathcal{O}(M_{\text{GUT}}^{2})$ at $M_{\text{GUT}}^{2}$ and $\mathcal{O}(m_{W}^{2})$ at $m_{W}^{2}$. There is no hierarchy problem.
@@
The fully renormalized mass function virtual ; Choi:2023:observables ; Choi:2025:Higgs shows that the scalar mass squared runs as the square of the energy scale $q^{2}$. This is not an additional assumption but a consequence of the standard quantum field theory tHooft:1971 ; tHooft:Veltman:1972 , once the renormalization is carried out completely. Therefore it is $\mathcal{O}(M_{\text{GUT}}^{2})$ at $M_{\text{GUT}}^{2}$ and $\mathcal{O}(m_{W}^{2})$ at $m_{W}^{2}$—the natural scaling of an unprotected quantity. In a grand unified theory, an order-one boundary condition at the unification scale suffices to produce an electroweak-scale Higgs, replacing the traditional $10^{-28}$ cancellation.
@@
The mass function constructed in Ref. virtual defines the scalar mass at every external momentum $q$. On-shell renormalization yields Choi:2023:observables ; Choi:2025:Higgs
@@
with the dressed propagator $i/(q^{2}-m^{2}(q^{2}))$ and pole mass $m_{h}$ fixed by $m^{2}(m_{h}^{2})=m_{h}^{2}$. [^1] Above particle-production thresholds, $\Sigma$ acquires an imaginary part; $\Sigma_{\text{ren}}$ denotes $\text{Re}\,\Sigma_{\text{ren}}$ throughout
The mass function is finite Bogoliubov:Parasiuk:1957 ; Hepp:1966 ; Zimmermann:1969 , and regularization-independent Choi:2023:observables ; Choi:2024:decoupling . In particular, the $q^{2}$ dependence of $m^{2}(q^{2})$ is a physical effect that is present regardless of the renormalization scheme. The contribution from gauge bosons can be made gauge-parameter invariant by absorbing non-invariant terms to the vertices, thanks to the Ward–Takahashi identity. The result coincides with the Feynman gauge choice, which we assume from now on Cornwall:1982 ; Binosi:Papavassiliou:2009 ; Papavassiliou:1997fn ; Papavassiliou:1997pb ; virtual .
What runs is the mass function $m^{2}(q^{2})$, the full momentum-dependent dressing. The pole mass is a fixed number; it does not run. The power-law behavior at momenta far from the pole is a prediction; the pole mass itself is data.
@@
    This verifies the ’t Hooft account (1). The crucial difference is that the ’t Hooft estimate gives a correction at $M_{\text{GUT}}$, whereas the renormalized mass function $m^{2}(q^{2})$ is defined at every scale and is computed bottom-up from the measured pole mass. Using $m_{h}=125.08$ GeV ATLAS:2023 ; CMS:2025 , the complete one-loop correction to the Higgs mass parameter (see Apeendix) gives, at $\sqrt{q^{2}}=M_{\text{GUT}}=10^{16}$ GeV,
@@
    which is $\mathcal{O}(1)$. The mass function, measured in units of the scale at which it is probed, is order one at every scale.
@@
The 28 orders of magnitude between $m_{h}^{2}$ and $M_{\text{GUT}}^{2}$ are the natural scaling of a dimension-2 quantity that runs as $q^{2}$—not a hierarchy that requires explanation. The reason is nothing but the absence of symmetry protection of the scalar mass.
@@
In the $\overline{\text{MS}}$ scheme Bardeen:Buras:Duke:Muta:1978 with dimensional regularization, quadratic powers of the renormalization scale $\mu^{2}$ are never generated—they correspond to poles at $d=2$, which are invisible at $d=4-\epsilon$. The beta function for $\bar{m}^{2}(\mu)$ is therefore purely logarithmic, and the scalar mass appears to run like a fermion mass. With no $\mu^{2}$ term in sight, there is nothing to trigger alarm—and the power-law $q^{2}$ dependence, which lives in the finite parts of the amplitude, goes unrecognized. This does not mean that dimensional regularization is at fault—the mass function $m^{2}(q^{2})$ is regularization-independent, as noted above—but that $\overline{\text{MS}}$ is a scheme in which the physical scale dependence remains implicit.
The validity of these results requires that the perturbation theory remain under control at all scales up to $M_{\text{GUT}}$ Degrassi:2012 ; Buttazzo:2013 . The one-loop correction $\Sigma_{\text{ren}}^{\text{SM}}/q^{2}\approx 0.86$ at $M_{\text{GUT}}$ is less than unity, so the propagator denominator $q^{2}-m^{2}(q^{2})=q^{2}(1-0.86)>0$ remains positive—no sign change, no ghost, no breakdown. The top quark alone gives +1.135+1.135, which exceeds unity, but the $W$ and $Z$ contributions (0.28-0.28 combined) bring the total below one; the full SM content is essential (see Appendix). This partial cancellation is not accidental: fermion loops carry a factor (1)(-1) from the closed Grassmann trace (spin-statistics theorem Pauli:1940 ), giving the top a positive contribution while the $W$ and $Z$ contribute with opposite sign. We have verified that the sign structure persists at two loops—each fermion loop carries its (1)(-1)—so the partial cancellation is a structural property of the SM’s spin content, not a one-loop coincidence. The two-loop correction is suppressed by a further factor of $m_{\text{EW}}^{2}/(16\pi^{2}v^{2})\log(M_{\text{GUT}}^{2}/m_{\text{EW}}^{2})$, giving $\Sigma^{(2)}/\Sigma^{(1)}\sim 0.1$$0.2$; this shifts the total from $0.86$ to roughly 0.90.91.01.0, which does not change the order-one character of $m^{2}(M_{\text{GUT}}^{2})/M_{\text{GUT}}^{2}$. That $r$ approaches unity at $M_{\text{GUT}}$ is expected: this is the boundary of the SM’s perturbative domain, where the GUT description takes over. Perturbation theory is controlled at every scale from the electroweak to the GUT.
@@
We now examine what this implies for a UV completion. The Standard Model must be completed at some high scale, whether by grand unification, string theory, or another framework. The scale $M_{\text{GUT}}\sim 10^{16}$ GeV indicated by gauge coupling unification Georgi:Quinn:Weinberg:1974 ; Langacker:Luo:1991 is representative of where such new physics is expected to appear; in $SU(5)$, the $X$ boson pole mass provides a more precise candidate for $M_{\text{GUT}}$. We take $SU(5)$ grand unification Georgi:Glashow:1974 as a concrete prototype; the specific parameters differ between UV completions, but the essential structure—an order-one boundary condition at the high scale, with the SM running below it—does not.
@@
The ratio $\Sigma_{\text{ren}}^{\text{SM}}/q^{2}$ grows logarithmically $\sim\log q^{2}$ and reaches unity near $10^{18}$ GeV, providing an upper bound on the SM’s domain of validity—in analogy with the Landau pole in the $U(1)_{Y}$ coupling, which independently requires new physics below that scale.
@@
where $\mu^{2}$ is the $\mathbf{5}$-Higgs mass parameter, $\lambda_{5}$ the coupling to the adjoint, $V$ the adjoint VEV, and $\Sigma^{\text{heavy}}_{\text{GUT}}$ contains loops from $X$, $Y$ bosons, the triplet, and the adjoint. This is the matching and the value $r$ provides a constraint on the GUT.
@@
an order-one requirement on $\mathcal{O}(1)$ parameters. In the traditional framework, the same constraint was an accuracy of $\mathcal{O}(10^{-28})$. The target has widened to a band of width $\sim 0.3$. The specific value $m_{h}=125$ GeV within the electroweak band is a precise constraint on the GUT—analogous to the constraint that the top Yukawa is $0.99$ rather than 0.50.5—but it is a constraint on $\mathcal{O}(1)$ parameters, not a fine-tuning.
@@
To sum up, the power-law running makes the scalar mass $m^{2}(q^{2})$ scale as $q^{2}$, so $m^{2}(q^{2})/q^{2}$ is $\mathcal{O}(1)$ at every scale. Any order-one GUT-scale boundary condition yields an electroweak-scale pole mass through the SM running. If the scalar mass ran logarithmically like a fermion mass, $m^{2}$ would barely change between scales, and $(125~\text{GeV})^{2}$ from $\mathcal{O}(M_{\text{GUT}}^{2})$ inputs would require $10^{-28}$—the very fine-tuning that the hierarchy problem describes.
@@
Appendix A Appendix: Explicit one-loop formulas
@@ Acknowledgements.
The author is grateful to Jaehoon Jeong, Taehyun Jung, Hyun-Min Lee and Matthew McCollough for discussions. This work is partly supported by the grant RS-2023-00277184 of the National Research Foundation of Korea. The author used Claude (Anthropic) for editing and discussion during the manuscript preparation. The final content was reviewed and approved by the author, who takes full responsibility for the work.